OpenStudy (anonymous):

what is the reference angle of 56 radians into degrees

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (ghazi):

56 radian =\[56 rad= \frac{ 180 }{ \pi }*56 degrees\]

OpenStudy (ghazi):

1 radian= 180/ pi degrees

OpenStudy (ghazi):

@emhart2012 clear?

OpenStudy (anonymous):

yeah but what is the reference angle?

OpenStudy (anonymous):

bc i get 3208.56365 for that and i need the reference angle

OpenStudy (ghazi):

did you convert 56 rad in to degrees

OpenStudy (anonymous):

yes and thus what i got but i need the reference angle of that

OpenStudy (anonymous):

i got it

OpenStudy (ghazi):

nice :)